### v1

# On the common index jump theorem and further developments

Abstract
In [LZ02], Long and Zhu established originally the common index jump theorem (CIJT) for symplectic paths in 2002, which was later generalized to its enhanced version (ECIJT) by Duan, Long and Wang in [DLW16] in 2016. Started from [GGM18] of 2018 and [GG20] of 2020, and finally in [CGG24] of 2024, a similar index theorem was obtained, i.e., Theorem 3.3 of [CGG24], which was called the index recurrence theorem there. In this short note, we give detailed proofs to show that the first 4 assertions dealing with index iterations in the total of 5 assertions in Theorem 3.3 of [CGG24] actually coincide completely with results in (ECIJT) of [DLW16].
1 Introduction
In the paper [LZ02] of Long and Zhu published in 2002, the common index jump theorem (CIJT for short below) was established, and since then this theorem has become a powerful tool in the study of many problems related to periodic solutions of Hamiltonian systems.
Note that in the paper [DLW16] of Duan, Long and Wang published in “Calculus of Variations” in 2016, CIJT was improved to the enhanced common index jump theorem (ECIJT for short below), i.e., Theorem 3.5 of [DLW16], with some more precise formula on index iterations. Then in recent [DLLW24] of Duan, Liu, Long and Wang in 2024, ECIJT was further generalized.
On the other hand, In [GGM18] of Ginzburg, Gürel and Macarini published in 2018, a so-called index recurrence theorem, Theorem 4.1 there, was obtained, and the authors wrote specially that “This result can be deduced from the so-called enhanced common index jump theorem due to Duan, Long and Wang [DLW16] (see also [Lo02], [LZ02])”. In [GG20] of Ginzburg and Gürel published in 2020, its Theorem 5.2 which was also called the index recurrence theorem, was obtained. In [CGG24] of Cineli, Ginzburg and Gürel appeared as arXiv:2410.13093v1 on October 16, 2024 and its revision arXiv:2410.13093v2 on July. 17, 2025, its Theorem 3.3 was obtained and was also called index recurrence theorem too, whose proof was a replenished version of and was more rigorous than that of Theorem 5.2 in [GG20]. Such a theorem became one of the main tools in the proof of the main multiplicity result on closed Reeb flow orbits in [CGG24].
Note that the Theorem 3.3 in [CGG24] contains total of 5 assertions (IR1)-(IR5), and it needs to be pointed out that the first 4 assertions there coincide completely with the identities (3.20)-(3.22) of Theorem 3.5 (ECIJT) and estimates (3.42) and (3.43) in Remark 3.6 in [DLW16]. Note also that the proofs of (IR1)-(IR4) in [CGG24] occupied the major part in the proofs of Theorem 3.3 in [CGG24]. On the other hand because all the assertions in Theorem 5.2 in [GG20] are covered by these (IR1)-(IR4) in [CGG24], and thus assertions of Theorem 5.2 in [GG20] coincide also with those proved in [DLW16] published in 2016 already too.
Because in [CGG24], the authors used the Arnold-Givental normal forms of real quadratic forms established in [AG], and in [DLW16] the authors used the normal forms and basic normal forms of real symplectic matrices established in [LD00], [LH99] and [Lo02] (cf. Subsections 1.4-1.9 on pp.16-47 and Chapter 8 in [Lo02] for details), the coincidences between above mentioned (IR1)-(IR4) of Theorem 3.3 in [CGG24] and results already proved in [DLW16] are not that obvious for usual readers. This short paper is devoted to give detailed proofs for such coincidences, which are contained specially in the 4 Propositions below.
In the following, when the page numbers of [CGG24] appear we are using those in arXiv:2410.13093v2 on July. 17, 2025.
2 Theorem 3.3 and Remark 3.6 of [DLW16] as well as Theorem 3.3 of [CGG24]
Note that in Theorem 3.5 on p.145 of [DLW16], the authors proved the following theorem,
Theorem 3.5 in [DLW16]. (The enhanced common index jump theorem for symplectic paths) Let ${\gamma}_{k}\in{\cal P}_{\tau_{k}}(2n)$ for $k=1,\ldots,q$ be a finite collection of symplectic paths. Let $M_{k}={\gamma}_{k}(\tau_{k})$. We extend ${\gamma}_{k}$ to $0,+\infty)$ by (3.9) of [DLW16] inductively. Suppose
$\hat{i}({\gamma}_{k},1)>0,\qquad\forall\;k=1,\ldots,q.\hskip 227.62204pt[DLW16,(3.17)].$
Let
$\check{m}\equiv\check{m}({\gamma}_{1},\ldots,{\gamma}_{q})=\min\{\check{M}_{k}\;|\;1\leq k\leq q\}.\hskip 170.71652pt[DLW16,(3.18)].$
where
$\check{M}_{k}=\left\{\matrix{\min\{k\in{\bf N}\;|\;k{\theta}\in 2\pi{\bf N},\;e^{\sqrt{-1}{\theta}}\in{\sigma}(M)\;{\rm with}\;{\theta}\in(0,2\pi)\cap\pi{\bf Q}\},&\;{\rm if}\;M\in{\rm Sp}_{vnu}(2n),\cr+\infty,&\;{\rm if}\;M\in{\rm Sp}_{cnu}(2n),\cr}\right$.
and
${\rm Sp}_{cnu}(2n)$ $=$ $\{M\in{\rm Sp}(2n)\;|\;\dim(M^{m}-I)=\dim(M-I),\;\forall\;m\in{\bf N}\},$
${\rm Sp}_{vnu}(2n)$ $=$ ${\rm Sp}(2n)\setminus{\rm Sp}_{cnu}(2n).$
Then for every integer $\bar{m}\in{\bf N}$, there exist infinitely many $(q+1)$-tuples $(N,m_{1},\ldots,m_{q})\in{\bf N}^{q+1}$ such that
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},1),\;\forall\;1\leq k\leq q,\;1\leq m<\check{m},\hskip 28.45274pt[DLW16,(3.19)],$
and the following hold for all $1\leq k\leq q$ and $1\leq m\leq\bar{m}$,
$\nu({\gamma}_{k},2m_{k}-m)=\nu({\gamma}_{k},2m_{k}+m)=\nu({\gamma}_{k},m),\hskip 128.0374pt[DLW16,(3.20)],$
$i({\gamma}_{k},2m_{k}+m)=2N+i({\gamma}_{k},m),\hskip 187.78836pt[DLW16,(3.21)],$
$i({\gamma}_{k},2m_{k}-m)=2N-i({\gamma}_{k},m)-2(S_{M_{k}}^{+}(1)+Q_{k}(m)),\hskip 85.35826pt[DLW16,(3.22)],$
$i({\gamma}_{k},2m_{k})=2N-(S_{M_{k}}^{+}(1)+C(M_{k})-2{\Delta}_{k}),\hskip 128.0374pt[DLW16,(3.23)],$
where as in [LZ02], we let
${\Delta}_{k}=\sum_{0<m_{k}{\theta}/\pi\}<{\delta}}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}),\hskip 213.39566pt[DLW16,(3.24)],$
and we define
$Q_{k}(m)=\sum_{{\theta}\in(0,2\pi),\;e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})\atop\{\frac{m_{k}{\theta}}{\pi}\}=\{\frac{m{\theta}}{2\pi}\}=0}S_{M_{k}}^{-}(e^{\sqrt{-1}{\theta}}).\qquad\hskip 142.26378pt[DLW16,(3.25)]$
Note that in (3.42) and (3.43) contained in Remark 3.6 on p.145 in [DLW16], the authors proved also the following two estimates.
Remark 3.6 (ii) in [DLW16]. By (4.10) and (4.40) in [LZ02] (cf. (11.1.10) and (11.2.14) of [Lo02]), we have
$m_{k}=\left(\left[\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right]+\chi_{k}\right)\bar{M},\qquad 1\leq k\leq q,\hskip 85.35826pt[DLW16,(3.42)]$
where $\chi_{k}=0$ or $1$ for $1\leq k\leq q$, $\bar{M}$ is a positive integer such that $\frac{\bar{M}{\theta}}{\pi}\in{\bf Z}$ whenever $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$ and $\frac{{\theta}}{\pi}\in{\bf Q}$ for some $1\leq k\leq q$, and we set $\bar{M}=1$ if no such eigenvalues exist. Furthermore, by (4.20) in Theorem 4.1 of [LZ02] (cf.(11.1.20) of [Lo02]), for any ${\epsilon}>0$, we can choose $N$ and $\{\chi_{k}\}_{1\leq k\leq q}$ such that
$\left|\left\{\frac{N}{\bar{M}\hat{i}({\gamma}_{k},1)}\right\}-\chi_{k}\right|<{\epsilon},\qquad 1\leq k\leq q.\hskip 113.81102pt[DLW16,(3.43)]$
In the above, $a]=\max\{k\in{\bf Z}\;|\;k\leq a\}$ and $\{a\}=a-[a]$ for all $a\in{\bf R}$.
On the other hand, in p.30 of [CGG24] put into arXiv on Oct. 16, 2024 and then on July.17,2025, under similar conditions the authors proved their Theorem 3.3 which contains five assertions (IR1)-(IR5). Here we are interested in the following first four assertions (IR1)-(IR4) on the index iteration theories of symplectic matrix paths started from the identity matrix.
Theorem 3.3 in [CGG24]. For any $\eta>0$ and any $\ell_{0}\in{\bf N}$, there exists a sequence $C_{l}\to\infty$ of real numbers and integer sequences $d_{il}\to\infty$ and $k_{ijl}\to\infty$ (as $l\to\infty$) such that for all $i$, $j$ and $l$, and all $\ell\in{\bf Z}$ in the range $1\leq|\ell|\leq\ell_{0}$, we have
$(IR1)\qquad|\hat{\mu}(\Phi_{ij}^{k_{ijl}})-d_{il}|<\eta,\hskip 256.0748pt(IR1-1)$
$\qquad\qquad{\it and\;in\;particular,}\;d_{il}=[\hat{\mu}(\Phi_{ij}^{k_{ijl}})],\;{\it and}\;$
$\qquad\qquad d_{il}-m\leq\mu_{-}(\Phi_{ij}^{k_{ijl}})\leq\mu_{+}(\Phi_{ij}^{k_{ijl}})\leq d_{il}+m;\hskip 147.95424pt(IR1-2)$
$(IR2)\qquad\mu_{\pm}(\Phi_{ij}^{k_{ijl}+\ell})=d_{il}+\mu_{\pm}(\Phi_{ij}^{\ell})\qquad{\it when}\quad 0<\ell\leq\ell_{0};$
$(IR3)\qquad\mu_{+}(\Phi_{ij}^{k_{ijl}-\ell})=d_{il}-\mu_{-}(\Phi_{ij}^{\ell})+({\beta}_{+}(\Phi_{ij}^{\ell})-{\beta}_{-}(\Phi_{ij}^{\ell})),$
$\qquad\qquad{\it where}\;0<\ell\leq\ell_{0}<k_{ijl}\;{\it and}\;{\beta}_{+}-{\beta}_{-}\leq m\;{\it with}\;{\beta}_{\pm}=0\;{\it when}\;\Phi^{\ell}_{ij}\;{\it is~non-degenerate};$
$(IR4)\qquad{\it Assertions\;(IR1)-(IR3)\;continue\;to\;hold\;with\;the\;same\;values\;of\;}d_{il}\;{\it and}\;k_{ijl}\;{\it when}$
$\qquad\qquad\Phi_{ij}\;{\it is\;replaced\;by\;its\;non-degenerate\;part\;}\Psi\;{\it with\;all\;invariants\;of\;}\Phi_{ij}\;{\it and\;its\;}$
$\qquad\qquad{\it iterates\;replaced\;by\;their\;counterparts\;for\;}\Psi,\;{\it i.e.,}\;m\;{\it replaced\;by\;}m^{\prime},\;\mu_{\pm}(\Phi_{ij}^{k_{ijl}})\;$
$\qquad\qquad{\it replaced\;by\;}\mu(\Phi_{ij}^{k_{ijl}}),\;{\beta}_{\pm}\;{\it set\;to\;be\;}0,\;{\it etc.},$
$(IR5)\qquad C_{l}-\eta<k_{ijl}a_{ij}<C_{l},\;{\it and}\;ka_{ij}<C_{l}-\eta\;{\it when}\;k<k_{ijl}\;{\it and}\;ka_{ijl}>C_{l}\;{\it when}\;k>k_{ijl}.$
Note that all of (IR1)-(IR4) are in fact claims on only one single symplectic path, and clearly (IR4) is a simple consequence of (IR2)-(IR3). Thus the authors of [CGG24] pointed out specially that their proofs for (IR1)-(IR3) can be simplified to the following three properties on a single symplectic path.
[…]
$d-m\leq\mu_{-}(\Phi^{k})\leq\mu_{+}(\Phi^{k})\leq d+m;\hskip 170.71652pt(IR1"-2)$
[…]
where $0<\ell\leq\ell_{0}<k$ and ${\beta}_{+}-{\beta}_{-}=b_{+}-b_{-}\leq m$ with ${\beta}_{\pm}=0$ when $\Phi^{\ell}$ is non-degenerate.
Next we explain the details of the comparisons of results contained in these 2 papers related to (IR1)-(IR4) in the following 4 sections.
3 The coincidence of ${\beta}_{-}({\gamma})$ in [CGG24] and $S_{{\gamma}(1)}^{-}(1)$ in [Lo02]
To get precise definitions of certain quantities used in [CGG24] and understand the proof of Theorem 3.3 in [CGG24], we make first the following concrete computations, and then prove the Proposition 1 below.
As in p.27 of [CGG24], using the normal form of symplectic matrices in Section 2.4 of [AG], the authors considered a totally degenerate matrix $A\in{\rm Sp}(2m)$. Then $A$ can be written into $A=\exp(JQ)$, where all eigenvalues of $JQ$ are zero. The quadratic form $Q$ can be symplectically decomposed into a sum of terms of the following three types according to Section 2.4 of [AG]:
(i) the identically zero quadratic form on ${\bf R}^{2\nu_{0}}$,
(ii) the quadratic form $Q_{0}=p_{1}q_{2}+p_{2}q_{3}+\cdots+p_{d-1}q_{d}$ in Darboux coordinates on ${\bf R}^{2d}$, where $d\geq 1$ is odd,
(iii) the quadratic forms $Q_{\pm}=\pm(Q_{0}+p_{d}^{2}/2)$ on ${\bf R}^{2d}$ for any $d\geq 1$.
Note that when the integer $d$ in the above (ii) is even, the quadratic form $Q_{0}$ can be split into a sum of sub-quadratic forms $Q_{i}$ defined on ${\bf R}^{2d_{i}}$ for $1\leq i\leq i_{0}$ for some integer $i_{0}\geq 2$ with each $d_{i}$ is an odd integer.
Step 1.1. On the invariants $b_{\ast}(Q)$ and ${\beta}_{\pm}(Q)$.
In [CGG24], for $\ast=0,\pm$, the authors defined
$b_{\ast}(Q)=\{\rm the\;number\;of\;Q_{\ast}\;{\rm appeared\;in\;the\;decomposition\;of\;}Q\}.$ (3.1)
Note that here $\nu_{0}(Q)$ is defined to be the number of identically zero quadratic forms in the decomposition of $Q$ on ${\bf R}^{2\nu_{0}}$.
Therefore for $A=\exp(JQ)$ there holds
$\nu(A)=\dim\ker(A-I)=2(b_{0}+\nu_{0})+b_{+}+b_{-},$ (3.2)
where $\nu(A)$ is the geometric multiplicity of the eigenvalue $1$ of $A$ as usual (cf. Definition 5.1.1 on p.111 of [Lo02]).
As shown below in our discussion, such a quadratic form $Q$ corresponds to a symplectic path $\Phi\in C([0,1],{\rm Sp}(2n))$. Then as described in p.27 of [CGG24], the authors decompose $\Phi$ into a sum of two paths $\Phi=\Phi_{0}\oplus\Psi$ such that $\Phi_{0}(1)\in{\rm Sp}(2n_{0})$ is totally degenerate and $\Psi(1)\in{\rm Sp}(2n_{1})$ is non-degenerate. In particular $n_{0}$ is the algebraic multiplicity of the eigenvalue $1$ of $\Phi(1)$ and $n_{0}+n_{1}=n$. Then the authors of [CGG24] defined first
$b_{\ast}(\Phi)=b_{\ast}(\Phi_{0}(1))\quad{\rm for}\;\;\ast=0,\pm\qquad{\rm and}\quad\nu_{0}(\Phi)=\nu_{0}(\Phi_{0}(1).$
Then they defined
${\beta}_{\pm}(\Phi)=\nu_{0}(\Phi)+b_{0}(\Phi)+b_{\pm}(\Phi).$ (3.3)
Step 1.2. On $b_{0}(\Phi)$.
We study the quadratic form $Q_{0}$ on ${\bf R}^{2d}$ with an odd integer $d\geq 1$ first.
Note that the values of the function $b_{0}$ introduced in p.27 of [CGG24] is defined by $b_{0}(Q)=\nu_{0}(Q_{0})$. From the definition of $Q_{0}$ in the above (ii), we obtain
$Q_{0}=\left(\matrix{O&A\cr A^{T}&O\cr}\right)_{2d\times 2d},\quad{\rm with\;}A=\left(\matrix{0&1&0&\cdots&0&0\cr 0&0&1&\cdots&0&0\cr 0&0&0&\cdots&0&0\cr\cdot&\cdot&\cdot&\cdots&\cdot&\cdot\cr 0&0&0&\cdots&1&0\cr 0&0&0&\cdots&0&1\cr 0&0&0&\cdots&0&0\cr}\right)_{d\times d},$ (3.4)
and $A^{T}$ is the transpose of $A$.
Consider the following initial value problem of the linear Hamiltonian system
$\dot{{\gamma}}(t)=JQ{\gamma}(t)\qquad\forall\;t\in[0,1],$ (3.5)
${\gamma}(0)=I.$ (3.6)
Here and below, according to Examples 2) in [AG,p.5], the standard symplectic matrix on the coordinate space ${\bf R}^{2n}$ is given by $J=\left(\matrix{0&-I_{n}\cr I_{n}&0\cr}\right)$ where $I_{n}$ is the $n\times n$ identity matrix on ${\bf R}^{n}$. According to the Corollary in p.5 of [AG], the Darboux coordinates is written as $(p_{1},\ldots,p_{d},q_{1},\ldots,q_{d})\in{\bf R}^{2n}$. Note that these choices coincides with those chosen in [Lo02] which are related to the orientation in the discussions here, and are important in the computations blow.
Then from (3.4)-(3.6) we obtain
${\gamma}(t)=\left(\matrix{B^{\ast}(t)&O\cr O&B(t)\cr}\right)_{2d\times 2d},\qquad\forall\;t\in[0,1],$ (3.7)
where
$B(t)=\left(\matrix{1&t&t^{2}/2&t^{3}/(3!)&\cdots&\cdots&t^{d-2}/((d-2)!)&t^{d-1}/((d-1)!)\cr 0&1&t&t^{2}/2&\cdots&\cdots&t^{d-3}/((d-3)!)&t^{d-2}/((d-2)!)\cr 0&0&1&t&\cdots&\cdots&t^{d-4}/((d-4)!)&t^{d-3}/((d-3)!)\cr 0&0&0&1&\cdots&\cdots&t^{d-5}/((d-5)!)&t^{d-4}/((d-4)!)\cr\cdot&\cdot&\cdot&\cdots&\cdots&\cdot&\cdot\cr 0&0&0&0&\cdots&\cdots&t&t^{2}/2\cr 0&0&0&0&\cdots&\cdots&1&t\cr 0&0&0&0&\cdots&\cdots&0&1\cr}\right)_{d\times d},$
and
$B^{\ast}(t)=\left(\matrix{1&0&0&0&\cdots&\cdots&0&0\cr(-t)&1&0&0&\cdots&\cdots&0&0\cr(-t)^{2}/2&(-t)&1&0&\cdots&\cdots&0&0\cr(-t)^{3}/3!&(-t)^{2}/2&(-t)&1&\cdots&\cdots&0&0\cr\cdot&\cdot&\cdot&\cdot&\cdots&\cdots&\cdot&\cdot\cr(-t)^{d-3}/((d-3)!)&(-t)^{d-4}/((d-4)!)&(-t)^{d-5}/((d-5)!)&(-t)^{d-6}/((d-6)!)&\cdots&\cdots&0&0\cr(-t)^{d-2}/((d-2)!)&(-t)^{d-3}/((d-3)!)&(-t)^{d-4}/((d-4)!)&(-t)^{d-5}/((d-5)!)&\cdots&\cdots&1&0\cr(-t)^{d-1}/((d-1)!)&(-t)^{d-2}/((d-2)!)&(-t)^{d-3}/((d-3)!)&(-t)^{d-4}/((d-4)!)&\cdots&\cdots&(-t)&1\cr}\right)_{d\times d}.$
Here we have
$\det({\lambda}I_{2d}-{\gamma}(t))=({\lambda}-1)^{2d},\qquad\forall\;t\in[0,1].$ (3.8)
Thus ${\sigma}({\gamma}(t))=\{1,1,\ldots,1\}$, i.e., $1$ is an eigenvalue of ${\gamma}(t)$ with algebraic multiplicity $2d$ for every $t\in[0,1]$.
By the study in Section 1.4 on pp.16-24 and Theorem 1.8.10 on p.41 of [Lo02], the matrix ${\gamma}(1)$ can be connected to a decomposition of basic normal forms in ${\Omega}_{1}^{0}({\gamma}(1))$, where the homotopy set ${\Omega}_{1}(M)$ respect to the eigenvalue $1$ is defined by
${\Omega}_{1}(M)=\{N\in{\rm Sp}(2d)\;|\;\nu_{1}(N)=\nu_{1}(M)\},$ (3.9)
and ${\Omega}_{1}^{0}(M)$ is the path connected component of ${\Omega}_{1}(M)$ containing $M$ defined in [Lo99b] and can be find in Definition 1.8.5 on p.38 of [Lo02]. Specially the above matrix ${\gamma}(t)$ with $t\in[0,1]$ is in the form of the matrix $N_{d}(1,0)$ defined in (1.4.1)-(1.4.5) with $b=0$ on pp.17-18 in [Lo02] which is written as
$N_{d}(1,0)=\left(\matrix{A_{d}(1)&O\cr O&C_{d}(1)\cr}\right)_{2d\times 2d}.$
Thus by Theorem 1.4.1 on p.18 of [Lo02], we obtain
$\nu({\gamma})=\dim\ker({\gamma}(1)-I)=2.$ (3.10)
where and below we write $\nu({\gamma})=\nu_{1}({\gamma})$ for notational simplicity, when the eigenvalue $1$ in consideration is clear enough.
Using the basic normal forms (defined in [Lo99a], [Lo00]) as in (1.8.15)-(1.8.17) on p.41 in [Lo02], by Theorem 1.8.10 on p.41 of [Lo02] again, it yields
$I_{2}{\diamond}H_{0}\in{\Omega}_{1}^{0}({\gamma}(1)),$ (3.11)
where $H_{0}\in{\rm Sp}(2(d-1))$ satisfies ${\sigma}(H_{0})\cap{\bf U}=\emptyset$.
In this case, authors of [CGG24] defined $b_{0}(Q)$ to be the number of $Q_{0}$ in the decomposition of $Q$. When $Q=Q_{0}$ given by (3.1) they defined
$b_{0}(Q_{0})=1.$ (3.12)
Consequently by the definition (3.3), it yields
$\beta_{\pm}(Q_{0})=b_{0}(Q_{0})=1.$ (3.13)
Using the splitting numbers $S^{\pm}_{M}({\lambda})$ defined in [Lo99b] (cf. Definition 9.1.4 on p.191 of [Lo02]), by the List 12 on p.198 of [Lo02] we have
$S_{{\gamma}(1)}^{\pm}(1)=S_{I_{2}}^{\pm}(1)=1.$ (3.14)
Thus, because both $S_{M}^{\pm}({\lambda})$ and $\beta_{-}(M)$ are additive in the decomposition of $M$ into sums of symplectic matrices, in this case we have
$S_{{\gamma}(1)}^{\pm}(1)=\beta_{\pm}({\gamma}).$ (3.15)
Step 1.3. On $b_{\ast}(\Phi)$ with $\ast=\pm$.
Next we study the quadratic form $Q_{\pm}$ on ${\bf R}^{2d}$ with an odd integer $d\geq 1$.
Because the quadratic forms $Q_{0}$ and $Q_{\pm}$ are operating on different subspaces of ${\bf R}^{2d}$, we can study them respectively, and then add them together to get the total quadratic form finally.
Subcase 1.3.1. The subcase $Q_{0}=0$.
In this subcase, by the definition in [AG] and [CGG24], the quadratic form $Q_{\pm}$ is defined on ${\bf R}^{2}$ as $Q_{{\epsilon}}$ with ${\epsilon}=1$ or 1-1 and the coordinates on ${\bf R}^{2}$ is given by $(p_{d},q_{d})$. In this case, from
$Q_{{\epsilon}}\left(\matrix{p_{d}\cr q_{d}\cr}\right)\cdot\left(\matrix{p_{d}\cr q_{d}\cr}\right)={\epsilon}\frac{p_{d}^{2}}{2},$
we obtain
$Q_{{\epsilon}}=\left(\matrix{{\epsilon}&0\cr 0&0\cr}\right).$
Then we consider the following initial value problem of the linear Hamiltonian system and we denote its fundamental solution by ${\gamma}_{{\epsilon}}(t)$ with ${\epsilon}=1$ or 1-1,
$\dot{{\gamma}}_{{\epsilon}}(t)=JQ_{{\epsilon}}{\gamma}_{{\epsilon}}(t),\qquad\forall\;t\in[0,1],$ (3.16)
${\gamma}_{{\epsilon}}(0)=I_{2}.$ (3.17)
Then solving this system directly, we obtain
${\gamma}_{{\epsilon}}(t)=\left(\matrix{1&0\cr{\epsilon}t&1\cr}\right),\qquad\forall\;t\in[0,1].$ (3.18)
Then
${\gamma}_{{\epsilon}}(1)=\left(\matrix{1&0\cr{\epsilon}&1\cr}\right)\in{\Omega}_{1}^{0}(N_{1}(1,-{\epsilon})),$ (3.19)
with $N_{1}(1,-{\epsilon})=\left(\matrix{1&-{\epsilon}\cr 0&1\cr}\right)$. Therefore by List 9.1.12 on p.198 of [Lo02], we obtain
$\nu({\gamma}_{-1})=\nu(N_{1}(1,1))=1,\qquad S_{{\gamma}_{-1}(1)}^{\pm}(1)=S_{N_{1}(1,1)}^{\pm}(1)=1,$ (3.20)
$\nu({\gamma}_{1})=\nu(N_{1}(1,-1))=1,\qquad S_{{\gamma}_{1}(1)}^{\pm}(1)=S_{N_{1}(1,-1)}^{\pm}(1)=0.$ (3.21)
Specially because ${\gamma}_{1}(1)\in{\Omega}_{0}(N_{1}(1,-1))$ and ${\gamma}_{-1}(1)\in{\Omega}_{0}(N_{1}(1,1))$ hold, it yields
${\beta}_{+}({\gamma}_{1})=1,\quad{\beta}_{-}({\gamma}_{1})=0,\quad{\rm and}\quad S_{{\gamma}_{1}(1)}^{\pm}(1)=S_{N_{1}(1,-1)}^{\pm}(1)=0,$ (3.22)
hold simultaneously, and
${\beta}_{+}({\gamma}_{-1})=0,\quad{\beta}_{-}({\gamma}_{-1})=1,\quad{\rm and}\quad S_{{\gamma}_{-1}(1)}^{\pm}(1)=S_{N_{1}(1,1)}^{\pm}(1)=1,$ (3.23)
hold simultaneously too. Thus, in this subcase we have always
$S_{{\gamma}_{\epsilon}(1)}^{-}(1)=\beta_{-}({\gamma}_{\epsilon})\qquad{\rm for}\quad{\epsilon}=\pm 1.$ (3.24)
Subcase 1.3.2. The subcase $Q_{0}\not=0$.
In this subcase, for ${\epsilon}=\pm 1$, by the definition in [AG] and [CGG24] the quadratic form $Q_{{\epsilon}}$ is given by $Q_{{\epsilon}}={\epsilon}(Q_{0}+p_{d}^{2}/2)$. Let ${\gamma}_{{\epsilon}}(t)$ be the fundamental solution of the initial value problem of the linear Hamiltonian system (3.16)-(3.17).
Then considering the above study on $Q_{0}$ on ${\bf R}^{2d}$ and $Q_{\pm}$ on ${\bf R}^{2}$, specially (3.3) and (3.10), by the study in Sections 1.8 and 1.9 on pp.36-47 in [Lo02], the basic normal form decomposition of ${\gamma}_{\pm 1}$ can be written as follows,
$I_{2}{\diamond}N_{1}(1,\mp 1){\diamond}H_{\mp 1}\in{\Omega}^{0}({\gamma}_{\pm 1}(1)),$ (3.25)
where $H_{\mp 1}\in{\rm Sp}(2(d-2))$ satisfies ${\sigma}(H_{\mp 1})\cap{\bf U}=\emptyset$.
Because the invariants in (3.3) and the splitting numbers $S_{M}^{\pm}({\lambda})$ in [Lo99b] and Definition 9.1.4 on p.191 of [Lo02] are symplectic additive, by (3.25) in the subcase 1.3.2 we obtain
$S_{{\gamma}_{{\epsilon}}(1)}^{-}(1)=\beta_{-}({\gamma}_{{\epsilon}}).$ (3.26)
Summarising the above studies, using the symplectic additivity of related invariants, we obtain
Proposition 1. For any ${\gamma}\in C([0,1],{\rm Sp}(2n))$, by the above proofs, the following holds always,
${\beta}_{-}({\gamma})=S_{{\gamma}(1)}^{-}(1).$ (3.27)
4 The (IR1) in Theorem 3.3 of [CGG24] is a direct consequence of Remark 3.6 of [DLW16]
In Section 3 of [CGG24], as usual the authors denote the set of all non-degenerate symplectic matrices in ${\rm Sp}(2m)$ with $m\in{\bf N}$ by ${\rm Sp}^{\ast}(2m)$. Then for a symplectic path $\Phi$ with end point $\Phi(1)\in{\rm Sp}^{\ast}(2m)$, the end point matrix $\Phi(1)$ can be connected to I-I in ${\rm Sp}^{\ast}(2m)$ by a path $\Psi$ lying entirely in ${\rm Sp}^{\ast}(2m)$. Let $\Phi^{\prime}$ be the concatenating path of $\Phi$ and $\Psi$. Then as usual the Conley-Zehnder index of $\Phi$ is defined by
$\mu(\Phi)=\hat{\mu}(\Phi^{\prime}),$ (4.1)
where $\hat{\mu}$ is defined to be the unique quasimorphism $\hat{\mu}:\widetilde{{\rm Sp}}(2m)\to{\bf R}$ which is continuous, homogeneous and satisfies the normalized condition:
$\hat{\mu}(\Phi_{0})=2\qquad{\rm for}\quad\Phi_{0}(t)=\exp(2\pi\sqrt{-1}t)\oplus I_{2m-2}.$
and $\widetilde{{\rm Sp}}(2m)$ is the universal covering of ${\rm Sp}(2m)$.
Then the authors of [CGG24] defined two indices for $\Phi\in\widetilde{{\rm Sp}}(2m)$ by
$\mu_{+}(\Phi):=\lim\sup_{\bar{\Phi}\to\Phi}\mu(\bar{\Phi}),\quad\mu_{-}(\Phi):=\lim\inf_{\bar{\Phi}\to\Phi}\mu(\bar{\Phi}),$ (4.2)
where in both cases the limit is taken over $\bar{\Phi}\in\widetilde{{\rm Sp}^{*}}(2m)$ converging to $\Phi$.
Our next result shows that the relations of the property (IR1”) with the indices defined in [Lo02] as well as [DLW16] are as follows,
Proposition 2. (i) The assertion (IR1”-1) in (IR1) are direct consequence of (3.42)-(3.43) in [DLW16].
(ii) The assertion (IR1”-2) in (IR1) was already proved in (1) of Theorem 1.1 of [LL00] (cf. $1^{o}$ of Theorem 10.1.2 on p.213 of [Lo02]).
Proof. Note first that using definitions of $\mu_{\pm}(\Phi)$ in (4.2), by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [Lo02], it yields immediately
$\mu_{-}(\Phi)=i(\Phi),$ (4.3)
$\mu_{+}(\Phi)=i(\Phi)+\nu(\Phi).$ (4.4)
Note that such identities were first introduced in [Lo90] of 1990 and rigorously proved in [Lo97] of 1997. Then they yield
$\hat{\mu}(\Phi)=\hat{i}(\Phi),\qquad\forall\;\Phi\in{\rm Sp}(2m).$ (4.5)
To prove (IR1”-1), letting $k$ and $d$ in (IR1”-1) be given by $k=2m_{i}$ and $d=2N$. Note first that using notations in [Lo02] and [DLW16], the left hand side of (IR1”-1) is given by
$|\hat{\mu}(\Phi_{i}^{k})-d|=|2m_{i}\hat{i}(\Phi_{i})-2N|.$
where $2m_{i}\hat{i}(\Phi_{i})=\hat{i}(\Phi_{i}^{2m_{i}})$ holds. Then by (3.42) and (3.43) of [DLW16], for any given $\eta>0$, we obtain
$|2m_{i}\hat{i}(\Phi_{i})-2N|$ $=$ $\left|2([\frac{N}{\bar{M}\hat{i}(\Phi_{i})}]+\chi_{i})\bar{M}\hat{i}(\Phi_{i})-2\frac{N}{\bar{M}\hat{i}(\Phi_{i})}\bar{M}\hat{i}(\Phi_{i})\right|$
$=$ $\left|-2\bar{M}\hat{i}(\Phi_{i})\left(\{\frac{N}{\bar{M}\hat{i}(\Phi_{i})}\}-\chi_{i}\right)\right|$
$<$ $2\bar{M}\hat{i}(\Phi_{i}){\epsilon}<\eta,$
when we take ${\epsilon}>0$ small enough. Therefore (IR1”-1) is proved.
Next using notations in [Lo02] and [DLW16], the integer $d$ in (IR1”-2) is given by $d=2N$ too. Therefore (IR1”-2) coincides precisely with (1) of Theorem 1.1 of [LL00] (cf. $1^{o}$ of Theorem 10.1.2 on p.213 of [Lo02]).
The proof is complete.
5 The coincidence of (IR2) and (IR3) in Theorem 3.3 of [CGG24] and (3.20)-(3.22) in Theorem 3.5 of [DLW16]
Now in the studies of the next two propositions, we prove that (IR2) and (IR3) of [CGG24] coincide completely with (3.20)-(3.22) of Theorem 3.5 in [DLW16]. In these proofs we use the notations in [DLW16] and [Lo02].
Proposition 3. The assertion (IR2”) and then (IR2) in [CGG24] coincides precisely with (3.20)-(3.21) in [DLW16].
Proof. Note first by Corollary 6.1.12 and Definition 6.1.13 on p.144 of [Lo02], it yields immediately (4.3)-(4.5). Therefore (IR2”) and then (IR2) in [CGG24] coincides precisely with (3.20)-(3.21) of [DLW16].
Next we have
Proposition 4. The assertion (IR3”) and then (IR3) in [CGG24] coincides precisely with (3.20)-(3.22) in [DLW16].
Proof. In order to use uniformly the notations in [DLW16] in the following proof, we have changed the notations ${\bf R}^{2m}$, ${\rm Sp}(2m)$, $\ell$ and $\Phi$ in (IR3”) of [CGG24] to ${\bf R}^{2n}$, ${\rm Sp}(2n)$, $m$ and $\gamma$} as those in (3.20) and (3.22) of [DLW16] respectively,
Note first that by the requirement of Theorem 3.3 of [CGG24] on the choice of $k$, the degree $D$ of ${\gamma}_{k}$ is a factor of $k$, and by the requirement (3.30) on $m_{k}$ of Theorem 3.5 of [DLW16], there holds $\frac{m_{k}{\theta}}{\pi}\in{\bf Z}$ whenever $\frac{{\theta}}{\pi}\in{\bf Q}\cap(0,2)$ and $e^{\sqrt{-1}{\theta}}\in{\sigma}(M_{k})$. These two requirements are precisely the same.
By (IR3”) in [CGG24] and (3.22) in [DLW16], to prove Proposition 4, it suffices to prove the following identity for every symplectic path in $C([0,1],{\rm Sp}(2n))$,
${\beta}_{+}({\gamma}_{k}^{m})-{\beta}_{-}({\gamma}_{k}^{m})=\nu({\gamma}_{k}^{2m_{k}-m})-2(S_{M_{k}}^{+}(1)+Q_{k}(m)),$ (5.1)
where $M_{k}={\gamma}_{k}(1)$, and $Q_{k}(m)$ is given by [DLW16,(3.25)].
By the above identities (3.2) and (3.3), we obtain
${\beta}_{+}({\gamma}_{k}^{m})+{\beta}_{-}({\gamma}_{k}^{m})=\nu({\gamma}_{k}^{m})=\nu({\gamma}_{k}^{2m_{k}-m}),$ (5.2)
where the last equality follows from [DLW16, (3.20)].
Then subtracting (5.2) from (5.1) yields
${\beta}_{-}({\gamma}_{k}^{m})=S_{M_{k}}^{+}(1)+Q_{k}(m)=S_{M_{k}}^{-}(1)+Q_{k}(m),$ (5.3)
where to get the last equality we have applied the fact $S_{M_{k}}^{+}(1)=S_{M_{k}}^{-}(1)$ from Lemma 9.1.6 on p.192 of [Lo02].
Then by the Bott-type formula for splitting numbers which was first proved in [LLZ02] (cf. Corollary 9.2.4 on p.201 in [Lo02]), we obtain
$S_{M_{k}}^{-}(1)+Q_{k}(m)=\sum_{\omega^{m}=1}S_{M_{k}}^{-}(\omega)=S_{M_{k}^{m}}^{-}(1),$ (5.4)
On the other hand, by Proposition 1 and the fact $M_{k}^{m}={\gamma}_{k}^{m}(1)$, we have
${\beta}_{-}({\gamma}_{k}^{m})=S_{M_{k}^{m}}^{-}(1).$
Together with (5.4), it yields a proof of (5.3) and then (5.1).
The proof of Proposition 4 is complete.